At time t=0 a car is moving 6m/s then smoothly accelerate so that the acceleration after t seconds is a a(t)=3t m/s



1
Expert's answer
2021-03-09T08:08:42-0500
"a(t)=\dfrac{dv}{dt}"

"dv=adt"

"\int dv=\int adt""v=\int adt"

Given "a=3t"


"v=\int(3t)dt"

"v(t)=\dfrac{3}{2}t^2+C_1"

If "v(0)=6"


"v(0)=\dfrac{3}{2}(0)^2+C_1=>C_1=6"

"v(t)=\dfrac{3}{2}t^2+6"


"v(t)=\dfrac{ds}{dt}"

"ds=vdt"

"\int ds=\int vdt""s=\int vdt""s=\int(\dfrac{3}{2}t^2+6)dt"

"s(t)=\dfrac{1}{2}t^3+6t+s(0)"

If "s(0)=0"


"s(t)=\dfrac{1}{2}t^3+6t"